Solve the equation. (If there are any unused answer boxes, enter NONE in the last boxes. If any real number is a solution, enter ANY in the answer boxes.)
|4x-5|-9=4
x = _-2__ (smaller value)
x = _4.5__ (larger value)

Answer by gonzo(654)

|4x-5|-9=4
becomes |4x-5| = 13.
this means that 4x-5 can be either + 13 or - 13 since the absolute value of 13 is 13 and the absolute value of -13 is also 13.
solving for 4x-5 = 13 yields x = 4.5
solving for 4x-5 = -13 yields x = -2
smaller value is -2.
bigger value is 4.5.